Jack drove 600 miles to a convention in Washington D.C. On the return trip he was able to increase his speed by 10mph and saved 3 hours of driving time. Find his rate for each direction.
Answer by lwsshak3(11628)

Jack drove 600 miles to a convention in Washington D.C. On the return trip he was able to increase his speed by 10mph and saved 3 hours of driving time. Find his rate for each direction.
**
let x=rate of speed to Washington D.C.
x+10=rate of speed for return trip
distance/rate of speed=travel time
600/x-600/(x+10)=3
LCD:x(x+10)
600(x+10)-600x=3x^2+30x
600x+6000-600x=3x^2+30x
3x^2+30x-6000=0
x^2+10x-2000=0
(x+40)(x-50)=0
x=-40(reject, x>0)
or
x=50
x+10=60
..
rate of speed to Washington D.C.=50 mph
rate of speed for return trip=60 mph